FOURIE TRANSFORM MICROWAVE SPECTRUM OF DIMETHYL ETHER DIMER

\underline{YOSHIO TATAMITANI}, BINGXIN LIU, JUN SHIMADA, T. OGATA, Department of Chemistry, Faculty of Science, Shizuoka University, Shizuoka, Japan 422-8529.

The microwave spectrum of dimethyl ether dimer (DME)_{2} has been observed using a Fourie transform microwave spectrometer at Shizuoka University from 5 to 18 GHz. The sample of 1% dimethyl ether in Ar was used to form the complex in a super sonic expansion. Sets of characteristic triplet lines with similar spectroscopic properties were found and assigned to be the J = 4←3, J = 5←4, and J = 6←5 transitions of (DME)_{2}. The microwave spectrum of the six isotopomers (DME)_{2}, (DME-d_{6})_{2}, (DME-^{13}C_{2})_{2}, (DME-d_{6})⋯(DME), (DME-^{13}C_{2})⋯(DME), (DME)⋯(DME-^{13}C_{2}) have been assigned, leading to the precise spectroscopic constants. From the rotational constants of these six isotopomers molecular structure of dimethyl ether dimer was determined, being found to have C_{s} symmetry and to be bound with three O⋯HC hydrogen bonds.The barrier height to internal rotation of the methyl group has been determined from the observed A-E splitting.